OpenStudy (anonymous):

The law of cosines is used to find the measures of a triangle when only the three sides are known, because the law of sines requires at least one angle measure. True or False?

OpenStudy (johnweldon1993):

This is true...because the law of sines says \[\frac{ \sin A }{ a } = \frac{ \sin B }{ b } = \frac{ \sin C }{ c }\] where A, B, and C are angle measurements...a,b and c are the side lengths.... If you do not know at least 1 angle...you cannot use this

OpenStudy (anonymous):

Thank you, thats what my book says but its stated in a twisty way

OpenStudy (johnweldon1993):

no problem!